A sample that weighs 109.75 g is a mixture of 40% helium atoms and 60% krypton atoms.

Part A

How many particles are present in the sample?

Express your answer using two significant figures.

Note: N= 7.1 x 10^24 is incorrect (what i got) and N= 1.1⋅10^25 is incorrect as well
Answer #1

Let there be N mole atom

Number of helium atom = 0.4*N

number of Krypton = 0.6*N

 

weight = 0.4*N*molar mass of helium + 0.6*N*molar mass of krypton

= 0.4*N*4 + 0.6*N*83.8

= 51.88*N

 

51.88*N = 109.75

N = 2.115 mol

= 2.115*6.022*10^23

=1.274*10^24 particle

 

Answer: 1.274*10^24 particle